The height of a projectile fired vertically into the air (neglecting air resistance) at an initial velocity of 160 feet per second is a function of time, x, and is given by the equation f(x) = 160x – 16x2 Find the highest point reached by the projective.
Answer by Alan3354(31530)

f(x) = -16x^2 + 160x
The vertex is on the line of symmetry.
The line of symmetry is x = -b/2a = -160/-32 = 5
f(5) = -400 + 800 = 400 feet